3л If cos x = - , and < x < , what is sin(x + )? Enter your answer as a fraction in simplest form, like this: 3/14
Answer #1

Cost = -4/5 TI CX <3I A 5 3 find sin(x + II) By Pythagarus theorem Ac²= A8²7 BC2 =) A8² - A C²-BC² =(5²-445 = 25-96=999 Hypot